# Euclid Quick Data Release (Q1) - Spectroscopic search, classification and analysis of ultracool dwarfs in the Deep Fields

### v1

@@ 3.4 Spectra for Žerjal’s catalog
We have confirmed and classified the brighter L- and T- type UCD candidates from M. Žerjal et al. (2025), beginning from the redder ones in the - color. To date, we have classified close to one hundred UCDs not included in Zhang’s compilation nor in the spectroscopic search, of which 7 are T dwarfs, 2 are late-L dwarfs, 40 are early L dwarfs, and 35 are late M dwarfs. This list includes some peculiar objects, such as E519971, E643591, E644720, E620204, and E524709. We are investigating their nature. Fig. 7 shows the spectra of the late L- and T-type UCDs already classified. All the new objects confirmed spectroscopically are listed in Table 8 (Appendix D).
3.5 Comparing the photometric and spectroscopic searches
There are six UCDs discovered by the spectroscopic search that were later found in the process of creating Žerjal’s catalog: E265716, E644720, E273015, E517518, E267056, and E271934. This overlap shows the good agreement in the selection criteria. There are also five UCDs discovered by the spectroscopic search algorithm that are not included in Žerjal’s catalog. Fig. 8 shows the - and - colors of the UCDs found by the two searches. The objects found exclusively in the spectroscopic search are:
@@ 3.5 Comparing the photometric and spectroscopic searches
This object is classified as a T5, is very faint in the visible ( = 27.6) and has colors - = 4 and - = 0. Even if its colors match the M. Žerjal et al. (2025) criteria it was not included in the photometric selection owing to the SNR_VIS flag.
@@
This object is classified as a T3 and the flags (det_quality_flag, flag_vis, flag_y, and flag_h) for the images are activated, so the color values have to be treated with caution.
@@
This object is classified as a T2, is faint in the visible ( = 25.4), with colors - = 3.8 and - = 0.6. It was not included in the photometric selection because of the ellipticity filtering flag.
@@
This object classified as a T1? owing to its peculiar spectrum. It is relatively bright in the visible ( = 23.2), with colors - = 3.0 and - = 0.5. It was filtered out from the photometric selection because of the mumax_minus_mag value.
From these objects we note that the main limitation is the quality of the images and derived data, given that the colors and filtering parameters are based on these data. It is worth to mention that E265716, the T4? object discussed in Sect. 3.3, is included in both the spectroscopic search and the photometric catalog. Its color values are: - = 3.9 and - = 0.7.
There are four objects selected by the photometric search that were not identified by the spectroscopic search:
@@
This object is an L8 dwarf that has P- and P- values below the minimum (0.6 and 0.5, respectively).
@@
This object is a peculiar T0 dwarf that also has P- and P- values below the minimum, owing in this case to the presence of artifacts or contamination around 1400 nm.
@@
This object is a T6 dwarf that was filtered out because of the presence of artifacts at different wavelengths producing NaN values in the search results.
@@ 5.2 Spectral indices
We also measured the NH3- spectral index that was redefined by E. L. Martín et al. (2021) anticipating Euclid low spectral resolution. This index is an indicator of the absorption of ammonia in late T and Y-type UCDs. We include earlier subtypes to investigate the dispersion of values obtained with Euclid spectra.
@@ 5.3 KI absorption doublet
NISP spectral range includes the region of the KI absorption doublet at 1243.6 / 1252.6 nm (air wavelength), and the spectral resolution is good enough to resolve the two lines. This potassium doublet is a known spectral feature that increases its depth with high gravity and therefore is related to age (R. O. Gray & C. Corbally, 2009). Fig. 12 shows an example of three M8 dwarfs from Q1 with different strengths in the potassium doublet. Given the different properties of the stellar populations in the EDFs and in the ERO-02 fields (E. Martín et al., 2024), we can compare old and young UCD populations. Fig. 13 shows two M7 objects as an example, E270809 from the EDF-N and E647131, a young object from the ERO-02 regions. The latter spectrum is taken from Dominguez-Tagle et al. (in prep.). The potassium doublet is strong in object E270809 whereas it is not present in E647131, a UCD from a young population. E647131 spectrum was extracted using our own code, also described in Dominguez-Tagle et al. (in prep.).
@@
We measured the pseudo equivalent width (pEW) of the KI absorption line 1252.6 nm in the spectra of 20 dwarfs that present the KI absorption doublet region not affected by artifacts. This list includes dwarfs from M7 to L4, and the results are presented in Table 4. To measure the pEW we deblend the two potassium lines and used a window of 20 nm for the pseudo continuum out of the line region. We do not estimate the pEW for the line at 1243.6 nm because it could be blended with an FeH feature at 1239 nm, which increases through the late-M dwarfs to early-L dwarfs (K. N. Allers & M. C. Liu, 2013). Fig. 14 shows the pEW values as a function of the spectral type. Object E269377 (WISEA J175730.71+674138.4) was observed with GTC/EMIR by J. Y. Zhang et al. (2024, see Sect. 2.1), and measurements from both spectra measured in the same manner are included in the figure. The difference between both measurements, 0.1 nm is considerably larger than the uncertainties. Other comparisons are needed to verify any systematics. Fig. 14 includes the estimations for the field UCDs reported by E. C. Martin et al. (2017) plotted as gray dots. These pEW values are lower than those for Q1 data (from M8 to L2), although they present a similar scatter and follow the same correlation with respect to the spectral type. Other authors show similar results (see e.g. K. N. Allers & M. C. Liu, 2013; N. Lodieu et al., 2018; A. J. Burgasser et al., 2025).
@@
The KI 1252.6 nm absorption line could also allow to estimate radial velocities, within the limits of the NISP low spectral resolution and calibration errors. The measurements of the wavelength accuracy reported by Y. Euclid Collaboration: Copin et al. (2025) indicate that 80% of the occurrences are below 0.25 nm. From this value, the minimum 3 reliable velocity estimate is 180 . Therefore, we do not aim to give an absolute value for the radial velocity, but to identify any object whose radial velocity deviates significantly from that of the rest of objects by comparing their values. Table 4 and Fig. 15 show the results. The figure is split into three panels, one per field. The objects in EDF-N and EDF-S present a dispersion of 0.4 and 0.5 nm RMS, respectively, which is larger than the typical wavelength accuracy reported by Y. Euclid Collaboration: Copin et al. (2025). The maximum difference between two objects in the EDF-S is 1.4 nm, that would be more than 330 . E269034, located in EDF-N, deviates by 1.1 nm (265 ) from the main group of objects in the same field. Further observations, either from ground-base telescopes or future visits of Euclid to these fields are needed to confirm these measurements. In the latter case, they will gradually improve the SNR, the wavelength calibration, and hence, the KI measurements.
@@ 6 Conclusions
We searched Q1 data for the UCD candidates from Zhang’s compilation and confirmed 224 objects. We selected 60 UCD benchmarks from this list to be the reference for the photometric selection criteria. The list includes 38 late M dwarfs, 19 L dwarfs, and 3 T dwarfs.
We searched for late-L and T-type UCDs directly in Q1 spectra and discovered 10 objects (8 T dwarfs and 2 late-L dwarfs).
Out of 5300 candidates in Žerjal’s catalog, we have (to date) confirmed close to one hundred UCDs that were not included in Zhang’s compilation, nor in the spectroscopic search. We compared the photometric and spectroscopic searches offering hints as to why some of them were not identified by the photometric selection.
In total, we discovered and classified 15 T dwarfs and 4 late-L dwarfs, combining the photometric and spectroscopic searches.
We created a preliminary list of Euclid UCD templates from M7 to T0, that can be used as a reference for the future search and classification of candidates.
@@ Appendix B UCD benchmarks
E581332 | T7 | -581332495491830038 | 03:52:32.0 | -49:10:58.8 | 17.8 |
E597913 | T7 | -597913643476826162 | 03:59:09.9 | -47:40:57.4 | 18.1 |
E607287 | T0 | -607287223470778395 | 04:02:54.9 | -47:04:40.2 | 18.9 |
E576081 | L4 | -576081983495392396 | 03:50:26.0 | -49:32:21.3 | 18.0 |
E638441 | L4 | -638441224479182813 | 04:15:22.6 | -47:55:05.8 | 17.7 |
E593750 | L4 | -593750792477935412 | 03:57:30.0 | -47:47:36.7 | 18.3 |
E594656 | L3 | -594656949486432578 | 03:57:51.8 | -48:38:35.7 | 18.3 |
E574238 | L3 | -574238742505125019 | 03:49:41.7 | -50:30:45.0 | 17.8 |
E603529 | L3 | -603529747469286537 | 04:01:24.7 | -46:55:43.2 | 18.1 |
Appendix C List of objects found by the spectroscopic search
@@ Appendix D Spectroscopically confirmed UCDs from Žerjal’s catalog
E271006 | T2 | 2710066793674540980 | 18:04:01.6 | +67:27:14.8 |
E643591 | T1? | -643591214457657250 | 04:17:26.2 | -45:45:56.6 |
E274809 | T0 | 2748094058670347269 | 18:19:14.3 | +67:02:05.0 |
E644720 | T0? | -644720877461587627 | 04:17:53.3 | -46:09:31.5 |
E620204 | T0? | -620204558458613530 | 04:08:04.9 | -45:51:40.9 |
E524709 | T0? | -524709228264023336 | 03:29:53.0 | -26:24:08.4 |
E517518 | L9 | -517518361295768184 | 03:27:00.4 | -29:34:36.5 |
Appendix E Effective temperatures estimations

### v2

@@ 3.4 Spectra for Žerjal’s catalog
We have confirmed and classified the brighter L- and T- type UCD candidates from Žerjal et al. (in prep.), beginning from the redder ones in the - color. To date, we have classified more than one hundred UCDs not included in Zhang’s compilation nor in the spectral index search, of which 4 are T dwarfs, 4 are late-L dwarfs, 56 are early L dwarfs, and 40 are late M dwarfs. This list includes some peculiar objects, such as E643591, E644720, and E524709, and E516938. We are investigating their nature. Fig. 7 shows the spectra of the late L- and T-type UCDs already classified. All the new objects confirmed spectroscopically are listed in Table 8 (Appendix D).
3.5 Comparing the photometric and spectral index searches
There are six late-L and T dwarfs discovered by the spectral index search that were later included Žerjal’s catalog: E265716, E644720, E273015, E517518, E267056, and E271934. This overlap shows the good agreement in the selection criteria. There are also five T dwarfs discovered by the spectral index search algorithm that are not included in Žerjal’s catalog. Fig. 8 shows the - and - colors of the UCDs found by the two searches. The objects found exclusively in the spectral index search are:
@@ 3.5 Comparing the photometric and spectral index searches
This object, classified as a T5, is very faint in the visible ( = 27.6) and has colors - = 4 and - = 0. Even if its colors match the Žerjal et al. (in prep.) criteria it was not included in the photometric selection owing to the SNR_VIS flag 222https://euclid.esac.esa.int/dr/q1/dpdd/.
@@
This object is classified as a T3p owing to its peculiar spectrum, probably due to a close very bright object. The flags (det_quality_flag, flag_vis, flag_y, and flag_h) for the images are activated, so the color values have to be treated with caution.
@@
This object, classified as a T2:, is faint in the visible ( = 25.4), with colors - = 3.8 and - = 0.6. It was not included in the photometric selection because of the ellipticity filtering flag.
@@
This object is classified as a T1: and is relatively bright in the visible ( = 23.2), with colors - = 3.0 and - = 0.5. It was filtered out from the photometric selection because of the mumax_minus_mag value.
From these objects we note that the main limitation is the quality of the images and derived data, given that the colors and filtering parameters are based on these data.
Two objects included in both the spectral index search and the photometric catalog are worth mentioning. E265716 is a possible unresolved binary owing to its peculiar spectrum that fits a T4 in the -band and T2 in the -band. Its color values are: - = 3.9 and - = 0.7. The second object, E271934 is a possible binary with a close companion that appears to be resolved in Euclid images. Its peculiar spectrum fits an L9 in the -band and L4 in the -band. Its color values are: - = 3.5 and - = 0.8. A full discussion is included in Muñoz Torres et al. (in prep.).
There are four objects selected by the photometric search that were not identified by the spectral index search:
@@
This object is an L8: dwarf that has P- and P- values below the minimum (0.6 and 0.5, respectively).
@@
This object is an L9: dwarf that also has P- and P- values below the minimum, owing to the presence of artifacts or contamination around 1400 nm.
@@
This object is a T5: dwarf that was filtered out because of the presence of artifacts at different wavelengths producing NaN values in the search results.
@@ 5.2 Spectral indices
We also measured the NH3- spectral index that was redefined by E. L. Mart´ın et al. (2021) anticipating Euclid low spectral resolution. This index is an indicator of the absorption of ammonia in late T and Y-type UCDs. We include earlier subtypes to investigate the dispersion of values obtained with Euclid spectra.
@@ 5.3 KI absorption doublet
NISP spectral range includes the region of the KI absorption doublet at 1243.6 / 1252.6 nm (air wavelength), and the spectral resolution is good enough to resolve the two lines. This potassium doublet is a known spectral feature that increases its depth with high gravity and therefore is related to age (R. O. Gray & C. Corbally, 2009). Fig. 12 shows an example of three M8 dwarfs from Q1 with different strengths in the potassium doublet. Given the different properties of the stellar populations in the EDFs and in the ERO-02 fields (E. L. Mart´ın et al., 2025), we can compare old and young UCD populations. Fig. 13 shows two M7 objects as an example, E661338 from the EDF-S and E647131, a young object from the ERO-02 regions. The latter spectrum is taken from Dominguez-Tagle et al. (in prep.). The potassium doublet is strong in object E270809 whereas it is not present in E647131, a UCD from a young population. E647131 spectrum was extracted using our own code, also described in Dominguez-Tagle et al. (in prep.).
@@
We measured the pseudo equivalent width (pEW) of the KI absorption line 1252.6 nm in the spectra of 20 dwarfs that present the KI absorption doublet region not affected by artifacts. This list includes dwarfs from M7 to L4, and the results are presented in Table 4. To measure the pEW we deblend the two potassium lines and use a window of 20 nm for the pseudo continuum out of the line region. We do not estimate the pEW for the line at 1243.6 nm because it could be blended with an FeH feature at 1239 nm, which increases through the late-M dwarfs to early-L dwarfs (K. N. Allers & M. C. Liu, 2013). Fig. 14 shows the pEW values as a function of the spectral type.
Object E269377 (WISEA J175730.71+674138.4) was observed with GTC/EMIR () by J. Y. Zhang et al. (2024, see Sect. 2.1), and the values from both spectra measured in the same manner are included in the figure using symbols. The 0.1 nm difference between both measurements is likely due to the different spectral resolutions, as reported by M. A. Malkan et al. (2002). Three objects show higher pEWs: E545928 (M8), E529572 (M8), and E619358 (L2). E529572 and E619358 present strong KI 1252.6 nm lines which could indicate a slightly higher gravity or a difference in the spectral subtype. Additional observations are needed to confirm this possibility.
Fig. 14 includes the measurements for the field UCDs reported by E. C. Martin et al. (2017) plotted as gray dots. Other authors show similar results (see e.g. K. N. Allers & M. C. Liu, 2013; N. Lodieu et al., 2018; A. J. Burgasser et al., 2025). In contrast, our pEW values are systematically higher, although they present a similar scatter and follow the same correlation with respect to the spectral type. In order to verify any systematic effect, we took the spectra observed by I. S. McLean et al. (2003) of the objects reported by E. C. Martin et al. (2017) and measured the KI 1252.6 nm pEW using the same method applied to Q1 data. The differences among the pEW values obtained with our method and those by E. C. Martin et al. (2017) were small, with a mean value of 0.06 nm. Then, following M. A. Malkan et al. (2002) procedure, we measured again the pEW after degrading the spectral resolution to match that of NISP (). These new values were systematically higher relative to the original ones. We computed a second-order polynomial fit between the values of the original and the degraded spectra. The polynomial coefficients are shown in the legend of the figure. We applied this relation to our measurements of Q1 spectra to derive the expected values for a similar resolution to that of E. C. Martin et al. (2017). The agreement between both sets of measurements is reasonable. The most plausible explanation is the blending of the two potassium lines (1243.6 and 1252.6 nm) owing to the low NISP spectral resolution. This effect needs to be considered in future pEW measurements from Euclid spectra.
@@
The KI 1252.6 nm absorption line could also allow to estimate radial velocities, within the limits of the NISP low spectral resolution and calibration errors. The measurements of the wavelength accuracy reported by Y. Euclid Collaboration: Copin et al. (2025) indicate that 80% of the occurrences are below 0.25 nm. From this value, the minimum 3 reliable velocity estimate is 180 . Therefore, we do not aim to give an absolute value for the radial velocity, but to identify any object whose radial velocity deviates significantly from that of the rest of objects by comparing their values. Table 4 and Fig. 15 show the results. The figure is split into three panels, one per field. The objects in EDF-N and EDF-S present a dispersion of 0.4 and 0.5 nm RMS, respectively, which is larger than the typical wavelength accuracy reported by Y. Euclid Collaboration: Copin et al. (2025). The maximum difference between two objects in the EDF-S is 1.4 nm, that would be more than 330 . E269034, located in EDF-N, also appears to deviate from the main group. However, a recent follow-up observation with GTC/EMIR shows only a very small shift (green symbol). The large difference between the two measurements exceeds the minimum 3 threshold for a reliable value, based on the available wavelength accuracy report. Therefore, the current measurements from NISP spectra should be interpreted with caution. Further observations, either from ground-base telescopes or future visits of Euclid to these fields are needed to confirm these measurements. In the latter case, they will gradually improve the SNR, the wavelength calibration, and hence, the KI measurements.
@@ 6 Conclusions
We searched Q1 data for the UCD candidates from Zhang’s compilation and confirmed the UCD nature of 224 objects. We selected 60 UCD benchmarks from this list to be the reference for the photometric selection criteria. The list includes 37 late M dwarfs, 21 L dwarfs, and 2 T dwarfs.
We also searched for late-L and T-type UCDs directly in Q1 spectra and discovered 11 objects (8 T dwarfs and 3 late-L dwarfs).
Out of 5300 candidates in Žerjal’s catalog, we have (to date) confirmed more than one hundred UCDs that were not included in Zhang’s compilation, nor in the spectral index search. We compared the photometric and spectral index searches offering hints as to why some of them were not identified by the photometric selection.
In total, we discovered and classified 12 T dwarfs and 7 late-L dwarfs, combining the photometric and spectral index searches.
We created a preliminary list of Euclid UCD templates from M7 to T1 , that can be used as a reference for the future search and classification of candidates.
@@ Appendix B UCD benchmarks
| E529572 | M8 | -529572577269202972 | 03:31:49.7 | -26:55:13.1 | 16.8 |
| E531137 | M8 | -531137796271166336 | 03:32:27.3 | -27:06:59.9 | 18.0 |
| E531382 | L1 | -531382284267016817 | 03:32:33.2 | -26:42:06.1 | 17.5 |
| E535151 | L1: | -535151413283919886 | 03:34:03.6 | -28:23:31.2 | 18.6 |
| E545928 | M8 | -545928778272623795 | 03:38:22.3 | -27:15:44.6 | 17.2 |
| E564140 | L2: | -564140829495758229 | 03:45:39.4 | -49:34:33.0 | 18.0 |
| E566608 | L1: | -566608621486213957 | 03:46:38.6 | -48:37:17.0 | 17.6 |
| E573377 | M8 | -573377256498994280 | 03:49:21.1 | -49:53:57.9 | 17.0 |
Appendix C List of objects found by the spectral index search
@@ Appendix D Spectroscopically confirmed UCDs from Žerjal’s catalog
| E520726 | L0 | -520726476283921213 | 03:28:17.4 | -28:23:31.6 |
Appendix E Effective temperature estimation